Hi everyone again, I know I already asked something similar to this already but something's not working out for me.
Four times the larger of 2 numbers decreased by 3 times the smaller is 152. Three times the larger increased by 5 times the smaller is 288. Find the numbers.
I started with x=larger number and y=smaller number
4x-3y=152 (1)
3x+5y=288 (2)
Now I'm stuck.

Answer by checkley75(3666)

THERE ARE MANY WAYS TO SOLVE THESE 2 EQUATIONS- ADD, SUBTRACT, ELIMINATION OR SUBSTITUTON.
HERE WE USE THE ELIMINATION MATHOD:
MULTIPLY THE FIRST EQUATION BY 5 & THE SECOND ONE BY 3 AND THEN ADD
20X-15Y=760
9X+15Y=864
---------------------------------
29X=1624
X=1624/29
X=56 ANSWER. NOW SUBSTITUTE 56 FOR X & SOLVE FOR Y
4*56-3Y=152
224-3Y=152
-3Y=152-224
-3Y=-72
Y=-72/-3
Y=24 ANSWER.